[Solution by Dr. Abubacker] The roots are $\frac13,\frac14$: $\frac{0.1}{P(B)}\cdot\frac{0.1}{P(A)}=\frac1{12}$ gives $P(A)P(B)=0.12$, and $0.1\left(\frac1{P(A)}+\frac1{P(B)}\right)=\frac7{12}$ gives $P(A)+P(B)=0.7$. So $P(\bar A\cup\bar B)=1-0.1=0.9$ and $P(\bar A\cap\bar B)=1-(0.7-0.1)=0.4$. Ratio $=\frac94$. Answer (4).